A man wants to know the velocity of current flow at pasig river. He used a motor boat with a known speed of 400 meter/minute. He clocked 5 minutes downstream and clocked 7 minutes upstream. Determine the distance covered. Determine the current flow.
1
Expert's answer
2019-11-20T09:56:11-0500

We have


"(v+u)t_1=(v-u)t_2"

"(400+u)5=(400-u)7"

"u=\frac{200}{3}\frac{m}{min}=\frac{200}{3}\frac{m}{60s}=1.1\frac{m}{s}"